the cost of 3 markers and 2 pencils is 1.80. the cost of 4 markers and 6 pencils is 2.90. Whhat is the cost of each item?
Answer by nerdybill(7384)

the cost of 3 markers and 2 pencils is 1.80. the cost of 4 markers and 6 pencils is 2.90. Whhat is the cost of each item?
.
Lem m = cost of a marker
and p = cost of a pencil
3m + 2p = 1.80
4m + 6p = 2.90
.
Multiply top equation by 3 and bottom equation by -1 and combine to get:
9m + 6p = 5.40
-4m - 6p = -2.90
-----------------
5m = 2.50
m = .50 (50 cents per marker)
.
Plug the value of m back into the top equation:
3m + 2p = 1.80
3(.50) + 2p = 1.80
1.50 + 2p = 1.80
2p = .30
p = .15 (15 cents per pencil)